# Causality for Complex Continuous-time Functional Longitudinal Studies with Dynamic Treatment Regimes

### v1

@@ 1.1.1 Longitudinal data formats
It is evident and important to realize that both RLD and ILD are special cases of FLD, as shown in the left figure of Figure 2. Therefore, methods designed for FLD can be applied to RLD and ILD, but typically not vice versa. Existing frameworks tailored for RLD or ILD cannot be straightforwardly applied to FLD. Intentional or unintentional usage of existing frameworks on FLD, for instance, pruning FLD into RLD, is prone to errors and hence might lead to unwanted causal conclusions, which can have serious vital consequences and result in a waste of healthcare resources. See Ying (2022, 2024) for a detailed discussion of the distinctions between the three types of longitudinal data.
1.1.2 Treatment regimes
@@
In addition to the complexity introduced by functional longitudinal data, the presence of time-varying treatments and the need to consider dynamic treatment regimes further complicate causal inference in this setting. A Treatment Regime (TR), also referred to as a strategy, plan, policy, or protocol, is a rule to assign treatment at each time of follow-up (Hernán and Robins, 2020). It is called a Static TR (STR) if it only depends on past treatment, while it is called a Dynamic TR (DTR) if it further depends on past covariates and outcomes. It is called deterministic if it is a determined function of past information, while it is called stochastic or random if it is random. For example, in a cross-sectional study, the average treatment effect focuses on a contrast between the averages of counterfactual outcomes if treatment were to be set to everyone or no one. This is static because it does not depend on covariates and deterministic because it is fixed. Suppose we define a rule that any patient were to receive treatment with 50% chance, then this is a stochastic STR. If, over age 65 were not to be treated while under age 65 were to be treated, then this is a deterministic DTR. If the patient under age 65 has a 50% chance of being treated, then this is a stochastic DTR. Apparently, deterministic TRs and STRs are special cases of stochastic TRs and DTRs, respectively.
@@
It is important to note that previous literature almost exclusively investigates prespecified DTRs, that is, DTRs that are known and independent of the actual TR, though in most cases they have abused the name DTR. Recent works have started to consider DTRs that depend on the actual TRs. For example, Haneuse and Rotnitzky (2013) considered “modified treatment policy” to answer what happens if, say, some continuous treatment dosage were to be added by some values. Robins, Hernán and SiEBERT (2004); Taubman et al. (2009) considered the “threshold intervention” to answer what happens if some continuous treatment like physical activity were to be set to a threshold if below but otherwise kept unchanged. Kennedy (2019) considered the so-called “incremental interventions” to shift the odds of receiving treatment and answered what happens if everyone’s odds of receiving treatment were, for instance, doubled compared to the actual odds. Young, Hernán and Robins (2014); Díaz and van der Laan (2012) both considered DTRs depending on actual TRs, where Díaz and van der Laan (2012) focused on “shift interventions”, for example, the effect of a policy that encourages people to exercise more, leading to a population where the distribution of physical activity is shifted according to certain health and socioeconomic variables. Therefore, distinguishing treatment regimes is not only a mathematical consideration but, more importantly, about what causal questions one is raising. In this paper, we investigate all DTRs (ADTRs), that is, we allow DTRs to depend on either the actual TR or not.
1.2 Objectives and Contributions
Functional longitudinal data (FLD) presents unique challenges for causal inference that are not adequately addressed by existing methods developed for regular or irregular longitudinal data. The continuous-time nature of FLD, coupled with the presence of an infinite number of potential treatment-confounder feedback loops, requires a novel theoretical framework that can account for these complexities. Existing methods often rely on the presence of a joint density for the treatment and confounder processes, which is not available in the FLD setting, and may make simplifying assumptions that do not hold for FLD. Furthermore, most existing methods for FLD focus on STRs, which may not be realistic in many real-world settings.
The primary objective of this paper is to develop a rigorous theoretical framework for causal inference in FLD under ADTRs. This framework aims to address the limitations of existing methods and provide a foundation for the development of practical estimation methods and algorithms. Our goal is not only to integrate and build upon the existing literature but also to extend the boundaries of our understanding by developing a corresponding semiparametric theory. Note that both the longitudinal data (FLD) and the treatment regimes (ADTRs) are as general as they can be (weakest to none assumptions on their formats).
To develop our framework, we draw inspiration from the concept of the Riemann integral. In the context of integration, one cannot simply evaluate the function point by point, as there are uncountably infinite points along the real number line. The Riemann integral bridges this gap between the finite and the infinite by partitioning the domain into smaller intervals and considering the function’s behavior within each interval. As the number of intervals increases and their width tends to zero, the Riemann sum converges to the true integral value.
Similarly, in the context of causal inference for FLD, we face the challenge of dealing with an uncountably infinite number of treatment-confounder feedback loops. Attempting to handle each time point individually, as one might do in the case of regular longitudinal data (RLD), becomes intractable. Moreover, without considering one point at a time, we lose access to a well-defined density function, which creates a tension between the existing theory based on densities and the theory needed for FLD.
Guided by this intuition, our key contributions are as follows:
  1. 1.
    We generalize the potential outcome framework for FLD under ADTR, which is surprisingly a nontrivial task.
  2. 2.
    We propose a novel set of assumptions tailored to the FLD setting, enabling the identification of causal effects under ADTRs. We construct a comprehensive identification framework for causal inference in FLD under ADTRs, including the development of g-computation, inverse probability weighting (IPW), and doubly robust formulas.
  3. 3.
    We develop the semiparametric theory, including the derivation of the efficient influence curve (EIC), providing a basis for optimal estimators and inference procedures in the future.
  4. 4.
    We discuss the implications of our results for practice and provide guidance on interpreting our key assumptions in real-world settings. We demonstrate the generality of our framework by showing how it subsumes existing methods for regular and irregular longitudinal data as special cases.

### v2

@@ 6.1 Causal Inference for Non-Functional Longitudinal Data
Current causal frameworks were designed to handle RLD (Greenland & Robins, 1986; Robins, 1986) or ILD (Lok, 2008; Johnson & Tsiatis, 2005; Røysland, 2011; Hu & Hogan, 2019; Rytgaard et al., 2022). It is evident to realize that both RLD and ILD are special cases of FLD, as shown in the left figure of Figure 1. Therefore, methods designed for FLD can be applied to RLD and ILD, but typically not vice versa. Existing frameworks tailored for RLD or ILD cannot be straightforwardly applied to FLD. Intentional or unintentional usage of existing frameworks on FLD, for instance, pruning FLD into RLD, is prone to errors and hence might lead to unwanted causal conclusions, which can have serious vital consequences and result in a waste of healthcare resources. See Ying (2022, 2024) for a detailed discussion of the distinctions between the three types of longitudinal data.
6.2 Causal Inference for Functional Data
Existing research on causal inference has explored the realm of functional treatments and covariates within observational studies, as noted in works by (Miao et al., 2020; Zhang et al., 2021; Tan et al., 2022). The data format these studies investigate is consistent with the framework of our analysis. Nonetheless, our work sets itself apart by emphasizing the temporal aspect inherent in longitudinal studies, in contrast to the primary focus on point exposure in the mentioned literature. Therefore, theoretically, by focusing on the same outcomes and setting aside the treatment-confounder feedback mechanism, our study could expand upon their findings regarding identification.
The only exceptions that investigated causal inference for functional longitudinal data are Ying (2022, 2024) and Sun & Crawford (2022). However, Ying (2022, 2024) only investigated static treatment regimes, without any semiparametric theory. On the other hand, Sun & Crawford (2022) imposed stochastic differential equations with stringent parametric assumptions, also under static treatment regimes. This situation highlights a significant gap in methodological advancements within the field.
6.3 Treatment Regimes
A Treatment Regime (TR), also referred to as a strategy, plan, policy, or protocol, is a rule to assign treatment at each time of follow-up (Hernán & Robins, 2020).
  • Static vs Dynamic: A TR is called Static (STR) if it only depends on past treatment, while it is called Dynamic (DTR) if it further depends on past covariates and outcomes.
  • Deterministic vs Stochastic: ATR is called deterministic if it is a determined function of past information, while it is called stochastic (or random) if it is random.
  • Prespecified vs Actual TR dependent:
@@ 6.3 Treatment Regimes
For example, in a cross-sectional study, the average treatment effect focuses on a contrast between the averages of counterfactual outcomes if treatment were to be set to everyone or no one. This is static because it does not depend on covariates and deterministic because it is fixed. Suppose we define a rule that any patient were to receive treatment with 50% chance, then this is a stochastic STR. If, over age 65 were not to be treated while under age 65 were to be treated, then this is a deterministic DTR. If the patient under age 65 has a 50% chance of being treated, then this is a stochastic DTR. Apparently, deterministic TRs and STRs are special cases of stochastic TRs and DTRs, respectively.
@@
It is important to note that previous literature almost exclusively investigates prespecified DTRs, that is, DTRs that are known and independent of the actual TR, though in most cases they have abused the name DTR. Recent works have started to consider DTRs that depend on the actual TRs. For example, Haneuse & Rotnitzky (2013) considered “modified treatment policy” to answer what happens if, say, some continuous treatment dosage were to be added by some values. Robins et al. (2004); Taubman et al. (2009) considered the “threshold intervention” to answer what happens if some continuous treatment like physical activity were to be set to a threshold if below but otherwise kept unchanged. Kennedy (2019) considered the so-called “incremental interventions” to shift the odds of receiving treatment and answered what happens if everyone’s odds of receiving treatment were, for instance, doubled compared to the actual odds. Young et al. (2014); Díaz & van der Laan (2012) both considered DTRs depending on actual TRs, where Díaz & van der Laan (2012) focused on “shift interventions”, for example, the effect of a policy that encourages people to exercise more, leading to a population where the distribution of physical activity is shifted according to certain health and socioeconomic variables. Therefore, distinguishing treatment regimes is not only a mathematical consideration but, more importantly, about what causal questions one is raising. In this paper, we investigate all DTRs (ADTRs), that is, we allow DTRs to depend on either the actual TR or not.
In Table 1, we position our paper among the literature.
Model-based methods
RLD
Robins (1998); Ying et al. (2023);
Vansteelandt & Joffe (2014)
ILD
Lok (2001); Johnson & Tsiatis (2005); Røysland (2011, 2012);
Hu & Hogan (2019); Yang (2022); Røysland et al. (2022)
FLD Singer (2008); Commenges & Gégout-Petit (2009); Sun & Crawford (2022)
Estimand-based methods for STRs
RLD
Rosenbaum & Rubin (1983);
Hernán et al. (2000, 2001, 2002)
ILD Rytgaard et al. (2022)
FLD Ying (2022, 2024)
Estimand-based methods for PDTRs
RLD
Murphy et al. (2001); Robins & Hernan (2008);
Young et al. (2011); Chakraborty & Murphy (2014)
ILD Rytgaard et al. (2022)
FLD This paper
Estimand-based methods for ADTRs
RLD
Kennedy (2019); Díaz, Hoffman & Hejazi (2023);
Díaz, Williams, Hoffman & Schenck (2023); Wen et al. (2023)
ILD This paper
FLD This paper
7 Discussion
7.1 Summary of the Paper